Your task is to find the cost of capital (WACC) for a company. The company has three sources of capital available. The marginal tax rate for the company is 35%

Debt: 2000 discount bonds with $1000 par value, with 4 years to maturity. Bonds currently offer 5% yield to bondholders.

Preferred stock:14 000 shares outstanding with $90 market price and 5% yield.

Common stock:100 000 shares outstanding with the book value of $20 but currently trading at P/B= 2.0. One may apply CAPM to estimate the cost of equity. Inputs for estimations are: risk-free rate 2.8%, market risk premium 6% and the stock beta is 1.20.

Find the cost of each financing source, capital structure weights for each source and the WACC.
Answer #1

Answer:

Debt:

Discount bond par value = $1,000

Yield = 5%

Years to maturity = 4

Current price = PV = FV / (1 + Interest rate) Number of periods

= $1,000 / (1 + 5%) 4

=$822.70

Current price of bond = $822.70

Market value of 2000 bonds = 2,000 * $822.70 = $1,645,400

Yield (given) = before tax cost of debt = 5%

Preferred stock:

Market value = Number of preference shares outstanding * Price per share = 14,000 * $90= $1,260,000

Market value = $1,260,000

Yield = Cost of preference share = 5%

Common stock:

Book value per share = $20

Currently trading at P/B= 2.0

Current market price = Book value * P/B = $20 * 2 = $40

Market value of equity = 100,000 * $40 = $4,000,000

Cost of equity = Risk free rate + Beta * Market risk premium = 2.8% + 1.20 * 6%= 10%

Market value of equity = $4,000,000

Cost of equity =10%

Capital structure weights:

Debt Preference share $1,260,000 Common Stock $4,000,000 Total Market value % weightage 23.83% 18.25% 57.93% 100.00% $1,645,400 $6,905,400

WACC

WACC = Cost of equity * Equity weightage + Cost of preference share * Preference share weightage +before tax cost of debt * (1 - Tax rate) * Debt weightage

= 10% * 57.93% + 5% * 18.25% + 5% * (1 - 35%) * 23.83%

WACC = 7.48%